OpenStudy (anonymous):

form a polynomial f(x) with real coefficients having the given degree and zeros degree 5; zeros 7; -i; -8+i enter the polynomial f(x)a=?

OpenStudy (campbell_st):

well you need the conjugate pairs for the complex roots f(x) = a(x -7)(x -i)(x + i)(x -(8+i))(x +(8+i)) hope it helps

OpenStudy (anonymous):

yes that helped then what is next

OpenStudy (campbell_st):

well why not distribute... (x -i)(x + i) = x^2 = -1 or x^2 + 1

OpenStudy (anonymous):

ok

OpenStudy (campbell_st):

so you have f(x) = a(x -7)(x^2 + 1)(x -(8+i))(x +(8 + i)) just keep distributing...

OpenStudy (anonymous):

yes i did that but now im stuck

OpenStudy (campbell_st):

ok... so the other quadratic factor is \[(x -(8 + 1))(x +(x +i)) = x^2 - (8 + i)^2\]

OpenStudy (anonymous):

ok so what will f(x)a=

OpenStudy (campbell_st):

well it should be f(x) =a(.......) becuase f(x) may be a multiply of the equation e.g. roots x = 2, 3 f(x) = a(x -2)(x -3) f(x) = a(x^2 -5x + 6) the value of a is found by substituting a point thats on the curve and then solving for a.

OpenStudy (anonymous):

ok thank you